OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

90

OpenStudy (lgbasallote):

let x = first integer obviously the consecutive integer would be x + 1 the third integer would be x+2 their sum is 267 so x + x + 1 + x + 2 = 267 3x + 3 = 267 3x = 267 - 3 3x = 264 solve for x

OpenStudy (anonymous):

No its x= 88

OpenStudy (anonymous):

Lets say X= largest integer: then x + (x-1) + (x-2) = 267 x= 90

OpenStudy (lgbasallote):

then you add 2 to get the largest integer

OpenStudy (anonymous):

or yeah, 90 works too.